In a triangle, the measure of the second angle is 40º less than the measure of the first angle. The measure of the third angle is twice the first. What are the measures of each of the angles in the triangle?
Answer by checkley71(8403)

LET X=THE FIRST ANGLE.
LET X+40=THE SECOND ANGLE.
LET 2X=THE THIRD ANGLE.
X+(X+40)+2X=180
4X+40=180
4X=180-40
4X=140
X=140/4
X=35 DEGREES FOR THE FIRST ANGLE.
35+40=75 FOR THE SECOND ANGLE.
2*35=70 FOR THE THIRD ANGLE.
35+75+70=180
180=180